Solve by completing the square.
x^2 +2x - 63 = 0
I got (7, -9) which doesn't appear to be correct. Can you help me?

Answer by edjones(8007)

x^2+2x-63=0
x^2+2x =63
x^2+2x+1=63+1
(x+1)^2=64
x+1=+-8
x=-1+-8
x=-9, x=7
.
ur correct
.
Ed